What annual rate of return is earned on a $3,400 investment when it grows to $7,300 in eighteen years?
Answer #1

Future Value = Present Value * (1+r)^n

7300 = 3400*(1+r)^18

(1+r)^18 = 7300/3400

= 2.14705882353

1+r = 2.14705882353^(1/18)

= 1.04336382407

r = 1.04336382407-1

= 0.04336382407

= 4.34%